OpenStudy (anonymous):

Susan Marciano invested part of her 40,000 bonus in a fund that paid an 11% profit and invested the rest in stock that surffered a 3% loss. Find the amount of each investment if her overall net profit was $1180.

OpenStudy (anonymous):

11% * x + -3% * (40,000-x) = $1180 2380 .11x -.03(40000-x) = 1180 .11x - 1200 + .03x =1180 .14x = 1180+1200 .14x = 2380 x= 2380/.14 = 17,000 so she invested $17,000 dollars in the investment that gained 11% and $23,000 in the other one

OpenStudy (anonymous):

disregard the 2380 on the second line...it's there by mistake